# The Complexity of HyperQPTL

### v2

@@ 4.1 Syntax and Semantics
In $\mathrm{Hyper\-QPTL}$ quantification over a proposition q is interpreted as labeling each trace by the same sequence $t_{\texttt{q}}$ of truth values for q, i.e., the assignment of truth values is uniform. However, one can also consider a non-uniform labeling by truth values for q. This results in the logic $\mathrm{Hyper\-QPTL^{+}}$.
@@ 4.2 Second-order HyperLTL
where p ranges over $\mathrm{AP}$, $\pi$ ranges over $\mathcal{V}_{1}$, $X$ ranges over $\mathcal{V}_{2}$, and $\mathop{\mathbf{X}\vphantom{a}}\nolimits$ (next) and $\operatorname{\mathbf{U}}$ (until) are temporal operators. Conjunction, implication, and equivalence are defined as usual, as are the temporal operators eventually ($\mathop{\mathbf{F}\vphantom{a}}\nolimits$) and always ($\mathop{\mathbf{G}\vphantom{a}}\nolimits$).
@@
For a variable assignment $\Pi$ and a position $i$ we define [^3] These are the standard semantics where second-order quantification ranges over arbitrary sets [1], not the closed-world semantics where second-order quantification only ranges over subsets of the model [8].
@@
  • $\Pi,i\models\psi_{1}\operatorname{\mathbf{U}}\psi_{2}$ if there is a $i^{\prime}\geq i$ such that $\Pi,i^{\prime}\models\psi_{2}$ and for all $i\leq i^{\prime\prime}<i^{\prime}$ we have $\Pi,i^{\prime\prime}\models\psi_{1}$,
@@ 4.3 HyperQPTL+ “is” Second-order HyperLTL
Let $\varphi$ be a $\mathrm{Hyper^{2}LTL}$ sentence. We assume without loss of generality that each (trace and set) variable is quantified at most once in $\varphi$. Further, we require that each set variable quantified in $\varphi$ is different from $X_{d}$ and $X_{a}$. These properties can always be achieved by renaming variables.
Our goal is to mimic quantification over sets of traces in $\mathrm{Hyper\-QPTL^{+}}$ via (non-uniform) quantification over propositions. Let $\texttt{p}_{1},\ldots,\texttt{p}_{n}$} be the propositions appearing in $\varphi$. Note that the truth values of other propositions can be ignored, as satisfaction of $\varphi$ does not depend on them. We will existentially quantify fresh propositions $\texttt{p}_{1}^{\mathit{all}},\ldots,\texttt{p}_{n}^{\mathit{all}}$ and then express in $\mathrm{Hyper\-QPTL^{+}}$ that all traces over these fresh propositions are available to mimick set quantification. Then, we use further propositional quantification in $\mathrm{Hyper\-QPTL^{+}}$ to mimic quantification over sets $X$ of traces by labeling the traces over $\texttt{p}_{1}^{\mathit{all}},\ldots,\texttt{p}_{n}^{\mathit{all}}$ with a bit indicating whether they are in $X$ or not. All other quantifiers, connectives, and temporal operators of $\mathrm{Hyper^{2}LTL}$ can directly be mimicked in $\mathrm{Hyper\-QPTL^{+}}$.
The formula $f(\varphi)$ we construct has the form
$f(\varphi)=\exists\texttt{p}_{1}^{\mathit{all}}.\ \ldots\exists\texttt{p}_{n}^{\mathit{all}}.\ (\theta_{\mathit{complete}}\wedge f^{\prime}(\varphi))$
with $\theta_{\mathit{complete}}$ and $f^{\prime}(\varphi)$ defined later. A nonempty set $T\subseteq(2^{\mathrm{AP}})^{\omega}$ of traces satisfies $f(\varphi)$ if and only if there is a set $T^{\prime}\subseteq(2^{\mathrm{AP}})^{\omega}$ of traces that satisfies both $\theta_{\mathit{complete}}\wedge f^{\prime}(\varphi)$ and $T=_{\mathrm{AP}\setminus\{\texttt{p}_{1}^{\mathit{all}},\ldots,\texttt{p}_{n}^{\mathit{all}}\}}T^{\prime}$.
@@
As a first step, we will express in $\mathrm{Hyper\-QPTL^{+}}$ that the $\{\texttt{p}_{1}^{\mathit{all}},\ldots,\texttt{p}_{n}^{\mathit{all}}\}$-projection of a set $T^{\prime}$ of traces contains all traces over $\{\texttt{p}_{1}^{\mathit{all}},\ldots,\texttt{p}_{n}^{\mathit{all}}\}$. To this end, consider the formula
@@
$\theta_{\mathit{complete}}=\forall\texttt{p}_{1}^{\mathit{temp}}.\ \ldots\forall\texttt{p}_{n}^{\mathit{temp}}.\ \forall\pi.\ \exists\pi^{\prime}.\ \bigwedge_{j=1}^{n}\mathop{\mathbf{G}\vphantom{a}}\nolimits((\texttt{p}_{j}^{\mathit{temp}})_{\pi}\leftrightarrow(\texttt{p}_{j}^{\mathit{all}})_{\pi^{\prime}})$
@@
with another set $\{\texttt{p}_{1}^{\mathit{temp}},\ldots,\texttt{p}_{n}^{\mathit{temp}}\}$ of fresh propositions used only in this subformula. Intuitively, it expresses that for each trace over $\{\texttt{p}_{1}^{\mathit{temp}},\ldots,\texttt{p}_{n}^{\mathit{temp}}\}$ the corresponding trace obtained by replacing each $\texttt{p}_{j}^{\mathit{temp}}$ by $\texttt{p}_{j}^{\mathit{all}}$ is in each model of $\theta_{\mathit{complete}}$, i.e., the $\{\texttt{p}_{1}^{\mathit{all}},\ldots,\texttt{p}_{n}^{\mathit{all}}\}$-projection of each model must indeed be the set of all traces over $\{\texttt{p}_{1}^{\mathit{all}},\ldots,\texttt{p}_{n}^{\mathit{all}}\}$.
@@
Now, let us explain how we mimic the quantification over subsets of traces in $\mathrm{Hyper\-QPTL^{+}}$: Similarly to the construction for the $\Sigma_{1}^{2}$-lower bound for $\mathrm{Hyper\-QPTL}$ satisfiability (see Lemma 1), we use propositional quantification over another fresh proposition m to mark each trace in the $\{\texttt{p}_{1}^{\mathit{all}},\ldots,\texttt{p}_{n}^{\mathit{all}}\}$-projection of a model (say at the first position) to indicate whether it is in the set or not. However, when quantifying m, we do not mark the $\{\texttt{p}_{1}^{\mathit{all}},\ldots,\texttt{p}_{n}^{\mathit{all}}\}$-projection of the model, but the model itself, i.e., we might label two different traces that have the same $\{\texttt{p}_{1}^{\mathit{all}},\ldots,\texttt{p}_{n}^{\mathit{all}}\}$-projection with different values of m. To rule this out, we require that the marking is consistent using the formula
@@
$\theta_{\mathit{cons}}^{j}=\forall\pi.\ (\mathop{\mathbf{X}\vphantom{a}}\nolimits\mathop{\mathbf{G}\vphantom{a}}\nolimits\neg(\texttt{m}_{j})_{\pi})\wedge\forall\pi^{\prime}.\ \bigwedge_{\ell=1}^{n}\Big(\mathop{\mathbf{G}\vphantom{a}}\nolimits\big((\texttt{p}_{\ell}^{\mathit{all}})_{\pi}\leftrightarrow(\texttt{p}_{\ell}^{\mathit{all}})_{\pi^{\prime}}\big)\Big)\rightarrow\Big((\texttt{m}_{j})_{\pi}\leftrightarrow(\texttt{m}_{j})_{\pi^{\prime}}\Big).$
@@
It expresses that only the first position can be marked (but must not) and that if two traces have the same $\{\texttt{p}_{1}^{\mathit{all}},\ldots,\texttt{p}_{n}^{\mathit{all}}\}$-projection then they also have the same marking. Now, let $X_{1},\ldots,X_{k}$ be the second-order variables quantified in $\varphi$. We employ a unique marker $\texttt{m}_{j}$ for each $j\in\{1,\ldots,k\}$.
@@
Note that traces in $T$ are still encoded by the “original” propositions $\texttt{p}_{1},\ldots,\texttt{p}_{n}$} while set quantification is mimicked by traces over the fresh propositions $\texttt{p}_{1}^{\mathit{all}},\ldots,\texttt{p}_{n}^{\mathit{all}}$}. Hence, when mimicking the binding of a trace from some $X_{j}$ or the distinguished variable $X_{a}$ to some $\pi$, we need to make sure that we use the fresh propositions and not the original ones. We do so by replacing each $(\texttt{p}_{j})_{\pi}$ by $(\texttt{p}_{j}^{\mathit{all}})_{\pi}$. Only when binding a trace from the distinguished variable $X_{d}$ (which ranges over the set $T$ we are evaluating $\varphi$ over) to some $\pi$ do we use the original propositions.
Thus, we define $f^{\prime}$ as
@@
An induction shows that we indeed have $T\models\varphi$ if and only if $T\models f(\varphi)$ for nonempty $T$: Over the structure of $\varphi$, one proves that for all nonempty sets $T$ of traces and for all variable assignments $\Pi$ of the free variables of a subformula $\psi$ that $T,\Pi,i\models\psi$ if and only if $T^{\prime},\Pi^{\prime},i\models f^{\prime}(\psi)$, where $T^{\prime}$ is obtained from $T$ by completely “overwriting” the propositions $\pi_{j}^{\mathit{all}}$ and assigning the marker propositions consistently according to $\Pi$, and where $\Pi^{\prime}$ is the restriction of $\Pi$ to trace variables. Again, we leave the tedious, but straightforward, details to the reader. ∎
@@
To mimic propositional quantification in $\mathrm{Hyper^{2}LTL}$, we use a dedicated second-order variable that explicitly stores the set of traces that the trace quantifiers in the $\mathrm{Hyper\-QPTL^{+}}$ formula range over. This set is initially equal to the set of traces the formula is evaluated over (i.e., the set $\Pi(X_{d})$ in the setting of $\mathrm{Hyper^{2}LTL}$), and is updated with each quantification over a proposition q. As this update from $T$ to $T^{\prime}$ has to satisfy $T=_{\mathrm{AP}\setminus\{\texttt{q}\}}T^{\prime}$, we need two set variables $X_{0}$ and $X_{1}$, one for the old value and one for the new value, to be able to compare these two. As the old value is no longer needed after the update, we can then reuse the variable. To keep track of which of the two variables $X_{0}$ and $X_{1}$ is currently used to store the set of traces we are evaluating trace quantifiers over, we use a flag $b\in\{0,1\}$, i.e., $X_{b}$ is the set variable storing the current value and $X_{1-b}$ is the variable storing the old value.
To enforce a correct update, we define
$\theta_{\texttt{q}}(X_{b},X_{1-b})={}$ ${}\forall\pi\in X_{b}.\ \exists\pi^{\prime}\in X_{1-b}.\ \bigwedge_{\texttt{p}\in\mathrm{AP}\setminus\{\texttt{q}\}}\mathop{\mathbf{G}\vphantom{a}}\nolimits(\texttt{p}_{\pi}\leftrightarrow\texttt{p}_{\pi^{\prime}})\wedge$
${}\forall\pi\in X_{1-b}.\ \exists\pi^{\prime}\in X_{b}.\ \bigwedge_{\texttt{p}\in\mathrm{AP}\setminus\{\texttt{q}\}}\mathop{\mathbf{G}\vphantom{a}}\nolimits(\texttt{p}_{\pi}\leftrightarrow\texttt{p}_{\pi^{\prime}}),$
@@
An induction shows that we indeed have $T\models\varphi$ if and only if $T\models f(\varphi)$: One shows that for all sets $T$ of traces and all variable valuations $\Pi$ of the free variables of a subformula $\psi$ that $T,\Pi,i\models\varphi$ if and only if $T,\Pi^{\prime},i\models f(\psi,b)$, where $\Pi^{\prime}$ is obtained from $\Pi$ by extending it with the appropriate assignment for the variable $X_{b}$, which must satisfy $T=_{\mathrm{AP}\setminus\{\texttt{q}_{1},\ldots,\texttt{q}_{k}\}}\Pi(X_{b})$, where the $\texttt{q}_{j}$ are the free propositional variables of $\psi$. One last time, we leave the tedious, but straightforward, details to the reader. ∎

### v3

@@ 4.1 Syntax and Semantics
In $\mathrm{Hyper\-QPTL}$, quantification over a proposition q is interpreted as labeling each trace by the same sequence $t_{\texttt{q}}$ of truth values for q, i.e., the assignment of truth values is uniform. However, one can also consider a non-uniform labeling by truth values for q. This results in the logic $\mathrm{Hyper\-QPTL^{+}}$.
@@ 4.2 Second-order HyperLTL
where p ranges over $\mathrm{AP}$, $\pi$ ranges over $\mathcal{V}_{1}$, $X$ ranges over $\mathcal{V}_{2}$, and $\mathop{\mathbf{X}\vphantom{a}}\nolimits$ (next) and $\operatorname{\mathbf{U}}$ (until) are temporal operators. Conjunction, implication, equivalence, and the temporal operators eventually and always are defined as usual.
@@
For a variable assignment $\Pi$ and a position $i$ we define
@@
  • $\Pi,i\models\psi_{1}\operatorname{\mathbf{U}}\psi_{2}$ if there is an $i^{\prime}\geq i$ such that $\Pi,i^{\prime}\models\psi_{2}$ and for all $i\leq i^{\prime\prime}<i^{\prime}$ we have $\Pi,i^{\prime\prime}\models\psi_{1}$,
@@
Note that these are the standard semantics where second-order quantification ranges over arbitrary sets [1], not the closed-world semantics where second-order quantification only ranges over subsets of the model [8].
@@ 4.3 HyperQPTL+ “is” Second-order HyperLTL
Let $\varphi$ be a $\mathrm{Hyper^{2}LTL}$ sentence. We assume w.l.o.g. that each (trace and set) variable is quantified at most once in $\varphi$. Further, we require that each set variable quantified in $\varphi$ is different from $X_{d}$ and $X_{a}$. These properties can always be achieved by renaming variables.
Our goal is to mimic quantification over sets of traces in $\mathrm{Hyper\-QPTL^{+}}$ via (non-uniform) quantification over propositions. Let $\texttt{p}_{1},\ldots,\texttt{p}_{n}$} be the propositions appearing in $\varphi$. Note that the truth values of other propositions can be ignored, as satisfaction of $\varphi$ does not depend on them. We existentially quantify fresh propositions $\texttt{p}_{1}^{\mathit{all}},\ldots,\texttt{p}_{n}^{\mathit{all}}$ and then express in $\mathrm{Hyper\-QPTL^{+}}$ that all traces over these fresh propositions are available to mimic set quantification. Then, we use further propositional quantification in $\mathrm{Hyper\-QPTL^{+}}$ to mimic quantification over sets $X$ of traces by labeling the traces over $\texttt{p}_{1}^{\mathit{all}},\ldots,\texttt{p}_{n}^{\mathit{all}}$ with a bit indicating whether they are in $X$ or not. All other quantifiers, connectives, and temporal operators of $\mathrm{Hyper^{2}LTL}$ can directly be mimicked in $\mathrm{Hyper\-QPTL^{+}}$.
In the following, we construct $\theta_{\mathit{cmplt}}$ and $f^{\prime}(\varphi)$ such that
$f(\varphi)=\exists\texttt{p}_{1}^{\mathit{all}}\ldots\exists\texttt{p}_{n}^{\mathit{all}}.\ (\theta_{\mathit{cmplt}}\wedge f^{\prime}(\varphi))$
which is satisfied by a nonempty set $T\subseteq(2^{\mathrm{AP}})^{\omega}$ of traces if and only if there is a set $T^{\prime}\subseteq(2^{\mathrm{AP}})^{\omega}$ of traces such that $T^{\prime}\models\theta_{\mathit{cmplt}}\wedge f^{\prime}(\varphi)$ and $T=_{\mathrm{AP}\setminus\{\texttt{p}_{1}^{\mathit{all}},\ldots,\texttt{p}_{n}^{\mathit{all}}\}}T^{\prime}$.
@@
As a first step, we need to express in $\mathrm{Hyper\-QPTL^{+}}$ that the $\{\texttt{p}_{1}^{\mathit{all}},\ldots,\texttt{p}_{n}^{\mathit{all}}\}$-projection of a set of traces contains all traces over $\{\texttt{p}_{1}^{\mathit{all}},\ldots,\texttt{p}_{n}^{\mathit{all}}\}$ using the formula
@@
$\theta_{\mathit{cmplt}}=\forall\texttt{p}_{1}^{\mathit{\!tmp}}\!\!\!\ldots\forall\texttt{p}_{n}^{\mathit{\!tmp}}.\ \forall\pi.\ \exists\pi^{\prime}.\!\bigwedge\nolimits_{j=1}^{n}\!\!\mathop{\mathbf{G}\vphantom{a}}\nolimits((\texttt{p}_{j}^{\mathit{\!tmp}})_{\pi}\!\leftrightarrow\!(\texttt{p}_{j}^{\mathit{all}})_{\pi^{\prime}})$
@@
with another set $\{\texttt{p}_{1}^{\mathit{\!tmp}},\ldots,\texttt{p}_{n}^{\mathit{\!tmp}}\}$ of fresh propositions used only in this subformula. Intuitively, it expresses that for each trace over $\{\texttt{p}_{1}^{\mathit{\!tmp}},\ldots,\texttt{p}_{n}^{\mathit{\!tmp}}\}$ the trace obtained by replacing each $\texttt{p}_{j}^{\mathit{\!tmp}}$ by $\texttt{p}_{j}^{\mathit{all}}$ is in each model of $\theta_{\mathit{cmplt}}$, i.e., the $\{\texttt{p}_{1}^{\mathit{all}},\ldots,\texttt{p}_{n}^{\mathit{all}}\}$-projection of each model must indeed be the set of all traces over $\{\texttt{p}_{1}^{\mathit{all}},\ldots,\texttt{p}_{n}^{\mathit{all}}\}$.
@@
Now, let us explain how we mimic the quantification over subsets of traces in $\mathrm{Hyper\-QPTL^{+}}$: Similarly to the construction for the $\Sigma_{1}^{2}$-lower bound for $\mathrm{Hyper\-QPTL}$ satisfiability (see Lemma 1), we use propositional quantification over another fresh proposition to mark each trace in the $\{\texttt{p}_{1}^{\mathit{all}},\ldots,\texttt{p}_{n}^{\mathit{all}}\}$-projection of a model (say at the first position) to indicate whether it is in the set or not. If $X_{1},\ldots,X_{k}$ are the second-order variables quantified in $\varphi$, then we employ a marker $\texttt{m}_{j}$ for each $j\in\{1,\ldots,k\}$.
However, we have to ensure that there are no two different traces that have the same $\{\texttt{p}_{1}^{\mathit{all}},\ldots,\texttt{p}_{n}^{\mathit{all}}\}$-projection, but different values of $\texttt{m}_{j}$. This can happen when quantifying $\texttt{m}_{j}$, as we do not mark the $\{\texttt{p}_{1}^{\mathit{all}},\ldots,\texttt{p}_{n}^{\mathit{all}}\}$-projection of the model, but the model itself. To rule this out, we require that the marking is consistent using the formula
@@
$\theta_{\mathit{cons}}^{j}=\forall\pi.\ (\mathop{\mathbf{X}\vphantom{a}}\nolimits\mathop{\mathbf{G}\vphantom{a}}\nolimits\neg(\texttt{m}_{j})_{\pi})\wedge\forall\pi^{\prime}.\bigwedge\nolimits_{\ell=1}^{n}\!\Big(\!\mathop{\mathbf{G}\vphantom{a}}\nolimits\big((\texttt{p}_{\ell}^{\mathit{all}})_{\pi}\leftrightarrow(\texttt{p}_{\ell}^{\mathit{all}})_{\pi^{\prime}}\big)\!\Big)\rightarrow\Big(\!(\texttt{m}_{j})_{\pi}\leftrightarrow(\texttt{m}_{j})_{\pi^{\prime}}\Big).$
@@
It expresses that only the first position can be marked (but must not) and that if two traces have the same $\{\texttt{p}_{1}^{\mathit{all}},\ldots,\texttt{p}_{n}^{\mathit{all}}\}$-projection they have the same marking.
@@
Note that traces in $T$ are still encoded by the “original” propositions $\texttt{p}_{1},\ldots,\texttt{p}_{n}$} while set quantification is mimicked by traces over the fresh propositions $\texttt{p}_{1}^{\mathit{all}},\ldots,\texttt{p}_{n}^{\mathit{all}}$}. Hence, when mimicking the binding of a trace from some $X_{j}$ or the distinguished variable $X_{a}$ to some $\pi$, we need to ensure that we use the fresh propositions and not the original ones. However, when mimicking the binding of a trace from $X_{d}$, we need to use the original propositions.
We do so by replacing each $(\texttt{p}_{j})_{\pi}$ by $(\texttt{p}_{j}^{\mathit{all}})_{\pi}$, unless when binding a trace from $X_{d}$ (which ranges over the set $T$ we are evaluating $\varphi$ over). Thus, we define $f^{\prime}$ as
@@
While $f^{\prime}(\varphi)$ is not necessarily in prenex normal form, it can easily be brought into prenex normal form, as there are no quantifiers under the scope of a temporal operator.
An induction shows that we indeed have $T\models\varphi$ if and only if $T\models f(\varphi)$ for nonempty $T$: Over the structure of $\varphi$, one proves that for all nonempty sets $T$ of traces and for all variable assignments $\Pi$ of the free variables of a subformula $\psi$ that $T,\Pi,i\models\psi$ if and only if $T^{\prime},\Pi^{\prime},i\models f^{\prime}(\psi)$, where $T^{\prime}$ is obtained from $T$ by completely “overwriting” the propositions $\texttt{p}_{j}^{\mathit{all}}$ and assigning the marker propositions consistently according to $\Pi$, and where $\Pi^{\prime}$ is the restriction of $\Pi$ to trace variables. Again, we leave the tedious, but straightforward, details to the reader. ∎
@@
To mimic propositional quantification in $\mathrm{Hyper^{2}LTL}$, we use a dedicated second-order variable that explicitly stores the set of traces that the trace quantifiers in the $\mathrm{Hyper\-QPTL^{+}}$ formula range over. This set is initially equal to the set of traces the formula is evaluated over (i.e., the set $\Pi(X_{d})$ in the setting of $\mathrm{Hyper^{2}LTL}$), and is updated with each quantification over a proposition q. As this update from $T$ to $T^{\prime}$ has to satisfy $T=_{\mathrm{AP}\setminus\{\texttt{q}\}}T^{\prime}$, we need two set variables $X_{0}$ and $X_{1}$, one for the old value and one for the new value, to be able to compare these two. As the old value is no longer needed after the update, we can then reuse the variable. To keep track of which of the two variables $X_{0}$ and $X_{1}$ is currently used to store the set of traces we are evaluating trace quantifiers over, we use a flag $b\in\{0,1\}$, i.e., $X_{b}$ is the set variable storing the current value and $X_{1-b}$ is the variable storing the old value. To enforce a correct update, we define the formula $\theta_{\texttt{q}}(X_{b},X_{1-b})$ as
${}\forall\pi\in X_{b}.\ \exists\pi^{\prime}\in X_{1-b}.\bigwedge\nolimits_{\texttt{p}\in\mathrm{AP}\setminus\{\texttt{q}\}}\mathop{\mathbf{G}\vphantom{a}}\nolimits(\texttt{p}_{\pi}\leftrightarrow\texttt{p}_{\pi^{\prime}})\wedge$
${}\forall\pi\in X_{1-b}.\ \exists\pi^{\prime}\in X_{b}.\bigwedge\nolimits_{\texttt{p}\in\mathrm{AP}\setminus\{\texttt{q}\}}\mathop{\mathbf{G}\vphantom{a}}\nolimits(\texttt{p}_{\pi}\leftrightarrow\texttt{p}_{\pi^{\prime}}),$
@@
An induction shows that we indeed have $T\models\varphi$ if and only if $T\models f(\varphi)$: One shows that for all sets $T$ of traces and all variable valuations $\Pi$ of the free variables of a subformula $\psi$ that $T,\Pi,i\models\psi$ if and only if $T,\Pi^{\prime},i\models f(\psi,b)$, where $\Pi^{\prime}$ is obtained from $\Pi$ by extending it with the appropriate assignment for the variable $X_{b}$, which must satisfy $T=_{\mathrm{AP}\setminus\{\texttt{q}_{1},\ldots,\texttt{q}_{k}\}}\Pi(X_{b})$, where the $\texttt{q}_{j}$ are the free propositional variables of $\psi$. One last time, we leave the tedious, but straightforward, details to the reader. ∎